The electric pot. at a point (x,0,0) is given by V =  [(1000/x) + (1500/x^2) + (500/x^2)] the the electric field at x = 1m is (in v/m)

Grade:11

1 Answers

RAGHVAN DAS
12 years ago

V=(1000/x)+(1500/x^2)+(500/x^2) at (x,0,0)---------(1)

E=-d(v)/dx

on differentiating equation 1 wrt to x we get

-[(1000/x^2)+2*((1500/x^3)+(500/x^3))]

E=-{-[(1000/x^2)+2*((1500/x^3)+(500/x^3))]}

now at x=1

E=1000+3000+1000=5000V/m

 

 

if your satisfied with the answer click on 'yes'